# Toward Stable Semi-Supervised Remote Sensing Segmentation via Co-Guidance and Co-Fusion

### v1

@@ III-B Explicit-Implicit Semantic Collaborative Guidance
For the CLIP-based student, we employ a concept-based prompt ensembling strategy to improve semantic precision. For each semantic category $k$, we curate a set of fine-grained concept descriptions $\mathcal{C}_{k}$ as defined in the dataset configurations. Each concept $c\in\mathcal{C}_{k}$ is formatted using a prompt template function $\mathcal{T}(c)$ and processed by the frozen CLIP text encoder $\mathcal{E}_{\text{txt}}$. We then aggregate these embeddings via averaging to obtain a representative class prototype $\mathbf{t}_{k}\in\mathbb{R}^{d_{1}}$ as follows:
$\mathbf{t}_{k}=\frac{1}{|\mathcal{C}_{k}|}\sum_{c\in\mathcal{C}_{k}}\mathcal{E}_{\text{txt}}\big(\mathcal{T}(c)\big).$ (5)
Subsequently, a lightweight projection layer $\phi(\cdot)$ maps these averaged embeddings into the decoder space to generate the final explicit query matrix $\mathbf{F}_{\text{clip}}^{t}\in\mathbb{R}^{K\times d_{1}}$:
$\mathbf{F}_{\text{clip}}^{t}=\phi([\mathbf{t}_{1},\dots,\mathbf{t}_{K}]),$ (6)
@@
In contrast, for the DINOv3-based student, we utilize learnable queries $\mathbf{Q}\in\mathbb{R}^{K\times d_{2}}$ to provide implicit class-level guidance. These queries are implemented as a trainable embedding layer and are optimized alongside the DINOv3 backbone and the segmentation decoder:
$\mathbf{Q}=[\mathbf{q}^{1},\dots,\mathbf{q}^{K}].$ (7)
@@
Let $\mathbf{F}_{\text{clip}}^{\text{global}}\in\mathbb{R}^{d_{1}\times h\times w}$ and $\mathbf{F}_{\text{dino}}^{\text{local}}\in\mathbb{R}^{d_{2}\times h\times w}$ denote the visual features from the CLIP-based and DINOv3-based students, respectively. Both students employ a homogeneous query-based decoder. The CLIP student produces the global prediction $\mathbf{M}_{\text{pred}}^{\text{global}}$ and the DINOv3 student produces the local prediction $\mathbf{M}_{\text{pred}}^{\text{local}}$. The segmentation logits are computed as:
$\mathbf{M}_{\text{pred}}^{\text{global}}=\mathrm{Sim}(\mathbf{F}_{\text{clip}}^{t},\mathbf{F}_{\text{clip}}^{\text{global}}),\quad\mathbf{M}_{\text{pred}}^{\text{local}}=\mathrm{Sim}(\mathbf{Q},\mathbf{F}_{\text{dino}}^{\text{local}}),$ (8)
where $\mathrm{Sim}(\cdot,\cdot)$ denotes a dot-product similarity implemented with an einsum operator. This formulation maintains a unified architectural structure while allowing the two students to leverage distinct semantic sources, namely explicit text embeddings versus implicit learnable queries.
@@ III-C Global-Local Feature Collaborative Fusion
Conventional co-training frameworks [chen2021semi, qiao2018deep, ke2019dual] typically rely on identical network architectures with different random initializations. Despite the divergent starting points, these homogeneous models often converge to similar error patterns due to confirmation bias. To mitigate this error accumulation, Co2S introduces a global-local feature collaborative fusion strategy. As visually evidenced in Fig. 3, the attention maps reveal distinct representational characteristics where CLIP produces diffuse activations rich in global contextual information while DINOv3 generates sharp responses highlighting local details. Building on this complementarity, our mechanism effectively fuses the global contextual information captured by CLIP [radford2021learning] with the local details produced by DINOv3 [simeoni2025dinov3]. This complementary fusion enables the model to generate segmentation maps with more accurate semantics and sharper, well-defined boundaries.
Given an unlabeled image $\mathbf{x}_{u}$, its weakly augmented view is fed into the two students to obtain the global logits $\mathbf{M}_{\text{pred}}^{\text{global}}$ and the local logits $\mathbf{M}_{\text{pred}}^{\text{local}}$. First, we compute the confidence maps $\mathbf{C}_{\text{pred}}^{\text{global}}\in\mathbb{R}^{1\times H\times W}$ and $\mathbf{C}_{\text{pred}}^{\text{local}}\in\mathbb{R}^{1\times H\times W}$ for the entire image:
$\mathbf{C}_{\text{pred}}^{\text{global}}=\max_{k}\big[\sigma(\mathbf{M}_{\text{pred}}^{\text{global}})\big]_{k},\quad\mathbf{C}_{\text{pred}}^{\text{local}}=\max_{k}\big[\sigma(\mathbf{M}_{\text{pred}}^{\text{local}})\big]_{k},$ (9)
where $\sigma(\cdot)$ denotes the softmax operator over the class dimension. To perform pixel-wise optimization, let $c_{\text{pred}}^{\text{global}},c_{\text{pred}}^{\text{local}}\in[0,1]$ denote the scalar confidence values at a specific spatial location from $\mathbf{C}_{\text{pred}}^{\text{global}}$ and $\mathbf{C}_{\text{pred}}^{\text{local}}$ respectively. The joint maximum confidence for this pixel is defined as:
@@
We then build interaction masks based on three confidence scenarios. When both students are confident with $c_{\text{pred}}^{\text{global}}\geq\tau$ and $c_{\text{pred}}^{\text{local}}\geq\tau$, the lower-confidence student learns from the higher-confidence one. If their confidence is equal, they engage in mutual learning. When only one student is confident with $C\geq\tau$ and the other is below $\tau$, the low-confidence student is supervised by the high-confidence peer. When both confidences are low with $c_{\text{pred}}^{\text{global}}<\tau$ and $c_{\text{pred}}^{\text{local}}<\tau$, the pixel is ignored and does not contribute to stabilization. Formally, the stability loss is defined as:
@@
where $\text{MSE}(\cdot,\cdot)$ represents the mean squared error calculation, $\mathbf{p}_{\ell},\mathbf{p}_{h}\in\mathbb{R}^{K}$ denote the probability vectors from the lower and higher-confidence students, $\tau$ is the confidence threshold, and $\lambda(t)$ is a ramp-up weight depending on the training step $t$. In the special case $c_{\text{pred}}^{\text{global}}=c_{\text{pred}}^{\text{local}}\geq\tau$, a symmetric mutual learning mechanism is applied.
@@ IV-A1 WHDLD
WHDLD [shao2020multilabel] comprises 4,940 optical images captured by GaoFen-1 and ZiYuan-3 satellites, each having a size of $\times256$ pixels and a 2-meter spatial resolution. It provides pixel-wise annotations for six categories: building, road, pavement, vegetation, bare soil, and water.
@@ IV-A2 LoveDA
LoveDA [wang2021loveda] provides 5,987 images of size $\times1,024$ pixels with 0.3-meter resolution. Spanning urban and rural scenes, it annotates seven categories: background, building, road, water, barren, forest, and agriculture.
@@ IV-A3 Potsdam
Potsdam [isprs-annals-I-3-293-2012] consists of 38 aerial orthophotos with 5-cm resolution and $\times6,000$ dimensions. It covers six classes: impervious surface, building, low vegetation, tree, car, and clutter. These images are cropped into 5,472 non-overlapping $\times512$ patches for training and evaluation.
@@ IV-A5 MER and MSL
FixMatch [sohn2020fixmatch]  [NeurIPS’20] 56.0 57.6 59.8 60.8
U2PL [wang2022semi]  [CVPR’22] 57.1 58.3 59.9 61.1
WSCL [lu2023weak]  [TGRS’23] 56.8 58.6 59.8 61.1
UniMatch [yang2023revisiting]  [CVPR’23] 57.4 58.8 60.4 61.5
DWL [huang2024decouple]  [ISPRS’24] 56.5 57.0 57.1 58.9
MUCA [wang2025semi]  [TGRS’25] 56.5 58.2 60.0 60.5
@@
TABLE II: Per-class IoUs(%) and mIoU(%) results on the LoveDA dataset under different labeled ratios. The best and second-best results are highlighted in bold and underlined, respectively.
1/40 Labeled Images
Method Background Building Road Water Barren Forest Agriculture mIoU
OnlySup 48.0 37.5 36.1 55.3 27.3 53.2 57.1 45.9
FixMatch [sohn2020fixmatch]  [NeurIPS’20] 50.3 52.1 53.3 69.9 28.7 56.4 62.3 53.3
U2PL [wang2022semi]  [CVPR’22] 47.4 47.5 50.9 69.8 30.0 56.5 63.8 52.3
WSCL [lu2023weak]  [TGRS’23] 52.9 52.1 62.8 68.1 31.2 57.7 63.8 54.1
UniMatch [yang2023revisiting]  [CVPR’23] 52.8 55.9 54.6 70.1 30.2 59.3 65.5 55.5
DWL [huang2024decouple]  [ISPRS’24] 50.4 57.8 57.7 75.7 36.4 58.5 65.1 57.4
MUCA [wang2025semi]  [TGRS’25] 49.2 50.8 51.0 73.5 33.9 58.9 64.7 54.6
Co2S (Ours) 51.6 59.7 58.3 76.5 35.9 61.5 64.2 58.2
1/16 Labeled Images
Method Background Building Road Water Barren Forest Agriculture mIoU
OnlySup 50.8 51.0 49.5 65.9 33.5 57.1 63.6 53.1
FixMatch [sohn2020fixmatch]  [NeurIPS’20] 55.5 58.4 55.4 72.2 37.4 59.3 66.9 57.9
U2PL [wang2022semi]  [CVPR’22] 53.5 58.3 56.6 72.6 38.0 59.9 67.3 58.0
WSCL [lu2023weak]  [TGRS’23] 53.8 55.8 55.2 71.6 37.5 61.3 66.6 57.4
UniMatch [yang2023revisiting]  [CVPR’23] 55.7 59.2 57.0 72.4 37.7 60.9 67.6 58.6
DWL [huang2024decouple]  [ISPRS’24] 52.7 60.6 59.1 73.7 36.0 62.9 68.2 59.0
MUCA [wang2025semi]  [TGRS’25] 52.6 54.1 54.5 75.0 34.1 61.2 65.0 56.7
Co2S (Ours) 54.0 62.2 59.0 76.3 40.0 63.6 67.6 60.4
1/8 Labeled Images
Method Background Building Road Water Barren Forest Agriculture mIoU
OnlySup 52.8 56.4 53.4 69.8 38.0 59.7 66.2 56.6
FixMatch [sohn2020fixmatch]  [NeurIPS’20] 55.9 60.9 58.8 73.1 40.6 62.3 69.2 60.1
U2PL [wang2022semi]  [CVPR’22] 54.8 60.9 59.4 72.0 40.3 61.6 68.4 59.6
WSCL [lu2023weak]  [TGRS’23] 56.6 60.0 58.5 72.9 38.6 61.1 68.0 59.4
UniMatch [yang2023revisiting]  [CVPR’23] 56.8 61.1 59.6 73.3 39.4 61.7 68.3 60.0
DWL [huang2024decouple]  [ISPRS’24] 56.2 61.5 60.0 78.9 47.1 66.6 69.4 62.8
MUCA [wang2025semi]  [TGRS’25] 55.5 57.4 58.2 77.3 38.3 65.4 69.2 60.2
Co2S (Ours) 57.2 63.8 60.4 79.3 42.3 66.7 69.2 62.7
1/4 Labeled Images
Method Background Building Road Water Barren Forest Agriculture mIoU
OnlySup 55.7 58.8 57.9 72.8 42.3 53.2 68.5 58.9
FixMatch [sohn2020fixmatch]  [NeurIPS’20] 56.9 61.9 60.6 73.8 43.3 64.2 69.8 61.5
U2PL [wang2022semi]  [CVPR’22] 57.4 61.5 60.6 74.6 42.9 64.3 69.2 61.5
WSCL [lu2023weak]  [TGRS’23] 57.6 61.3 59.3 74.3 43.4 63.2 68.6 61.1
UniMatch [yang2023revisiting]  [CVPR’23] 58.0 61.1 60.8 74.2 44.6 63.5 69.8 61.7
DWL [huang2024decouple]  [ISPRS’24] 56.5 61.8 60.6 79.3 44.5 68.8 72.4 63.4
MUCA [wang2025semi]  [TGRS’25] 55.9 61.8 61.1 78.0 44.8 68.3 71.6 63.1
Co2S (Ours) 57.0 63.4 61.0 79.1 48.3 67.8 71.2 64.0
@@
FixMatch [sohn2020fixmatch]  [NeurIPS’20] 69.4 71.5 73.5 75.7
U2PL [wang2022semi]  [CVPR’22] 67.2 71.3 73.8 75.9
WSCL [lu2023weak]  [TGRS’23] 69.9 72.3 73.9 75.4
UniMatch [yang2023revisiting]  [CVPR’23] 70.7 72.6 74.8 76.4
DWL [huang2024decouple]  [ISPRS’24] 74.2 78.3 79.8 80.3
MUCA [wang2025semi]  [TGRS’25] 66.9 71.5 76.0 79.2
@@
TABLE IV: Per-class IoUs(%) and mIoU(%) results on the MER dataset under different labeled ratios. The best and second-best results are highlighted in bold and underlined, respectively.
1/8 Labeled Images
Method Martian Soil Sands Gravel Bedrock Rocks Tracks Shadows Unknown Background mIoU
OnlySup 4.7 55.1 71.4 61.7 38.0 49.5 34.2 73.0 87.0 52.8
FixMatch [sohn2020fixmatch]  [NeurIPS’20] 2.0 53.6 73.5 65.8 40.0 54.8 31.5 72.4 90.4 53.8
U2PL [wang2022semi]  [CVPR’22] 5.2 51.8 73.2 70.5 45.7 54.6 26.2 73.8 93.2 54.9
WSCL [lu2023weak]  [TGRS’23] 3.5 54.8 69.9 62.8 36.0 56.9 26.1 76.0 80.9 51.9
UniMatch [yang2023revisiting]  [CVPR’23] 5.5 54.4 72.7 68.5 44.7 55.2 38.0 73.8 91.1 56.0
DWL [huang2024decouple]  [ISPRS’24] 5.9 51.0 74.9 67.2 41.4 61.9 28.3 69.7 90.4 54.5
MUCA [wang2025semi]  [TGRS’25] 5.3 56.0 72.6 60.5 36.1 52.9 34.5 67.7 87.8 52.6
Co2S (Ours) 3.0 56.3 73.1 67.9 37.0 66.3 36.5 79.7 91.3 56.8
1/4 Labeled Images
Method Martian Soil Sands Gravel Bedrock Rocks Tracks Shadows Unknown Background mIoU
OnlySup 21.6 56.3 74.4 66.3 38.0 52.3 22.1 74.5 92.2 55.3
FixMatch [sohn2020fixmatch]  [NeurIPS’20] 19.4 55.4 74.0 57.1 43.6 57.3 39.8 75.7 90.9 58.1
U2PL [wang2022semi]  [CVPR’22] 7.1 56.4 73.7 69.7 41.9 60.8 23.3 81.4 90.3 56.1
WSCL [lu2023weak]  [TGRS’23] 8.2 52.8 71.1 65.5 35.7 61.0 31.4 79.0 89.0 54.9
UniMatch [yang2023revisiting]  [CVPR’23] 19.7 54.6 74.3 69.3 44.5 54.9 37.4 77.4 90.5 58.1
DWL [huang2024decouple]  [ISPRS’24] 13.8 55.9 74.8 70.3 36.2 57.0 28.8 66.0 89.7 54.7
MUCA [wang2025semi]  [TGRS’25] 1.0 55.1 73.8 65.6 38.4 56.3 28.6 73.7 84.9 53.0
Co2S (Ours) 13.1 63.3 76.1 69.4 37.1 63.1 41.1 77.1 91.4 59.1
TABLE V: Per-class IoUs(%) and mIoU(%) results on the MSL dataset under different labeled ratios. The best and second-best results are highlighted in bold and underlined, respectively.
1/8 Labeled Images
Method Martian Soil Sands Gravel Bedrock Rocks Tracks Shadows Unknown Background mIoU
OnlySup 26.8 61.6 66.8 72.2 37.9 45.5 37.0 49.1 85.1 53.6
FixMatch [sohn2020fixmatch]  [NeurIPS’20] 31.0 65.4 69.6 75.0 40.4 49.6 39.8 54.2 86.5 56.8
U2PL [wang2022semi]  [CVPR’22] 26.7 64.7 70.3 73.3 39.8 43.1 40.3 53.1 87.7 55.5
WSCL [lu2023weak]  [TGRS’23] 25.8 65.0 68.3 73.9 38.8 69.6 39.6 57.0 86.0 58.2
UniMatch [yang2023revisiting]  [CVPR’23] 29.3 66.7 69.8 75.4 39.2 50.8 43.0 57.9 89.2 57.9
DWL [huang2024decouple]  [ISPRS’24] 33.5 65.9 71.2 74.7 44.7 48.8 44.8 56.3 90.9 59.0
MUCA [wang2025semi]  [TGRS’25] 15.4 63.9 67.0 71.5 40.8 33.4 41.0 55.9 82.9 52.4
Co2S (Ours) 36.7 68.4 71.7 76.0 42.6 58.9 44.1 60.9 89.3 60.9
1/4 Labeled Images
Method Martian Soil Sands Gravel Bedrock Rocks Tracks Shadows Unknown Background mIoU
OnlySup 32.2 65.0 69.0 75.1 43.6 57.6 44.0 57.9 86.6 59.0
FixMatch [sohn2020fixmatch]  [NeurIPS’20] 29.7 65.5 69.9 77.1 42.5 63.8 48.1 60.7 87.5 60.5
U2PL [wang2022semi]  [CVPR’22] 36.1 62.7 72.7 76.3 43.3 63.7 44.2 65.6 90.1 61.6
WSCL [lu2023weak]  [TGRS’23] 30.9 64.7 70.4 75.5 42.0 58.3 42.8 65.0 89.7 59.9
UniMatch [yang2023revisiting]  [CVPR’23] 32.2 67.7 72.6 77.0 42.5 71.8 49.1 61.9 90.8 62.8
DWL [huang2024decouple]  [ISPRS’24] 33.8 68.2 72.3 76.1 43.2 57.9 48.2 70.7 91.9 63.6
MUCA [wang2025semi]  [TGRS’25] 33.7 69.3 72.9 76.9 46.2 40.4 49.9 55.6 89.2 59.3
Co2S (Ours) 36.0 67.7 71.8 77.8 45.5 77.0 49.2 76.7 91.7 65.9
TABLE VI: Segmentation results (mIoU %) on the GID-15 dataset under different labeled ratios. The best and second-best results are highlighted in bold and underlined, respectively.
Method OnlySup FixMatch [sohn2020fixmatch] U2PL [wang2022semi] WSCL [lu2023weak] UniMatch [yang2023revisiting] DWL [huang2024decouple] MUCA [wang2025semi] Co2S (Ours)
@@
IV-B Implementation Details
The proposed Co2S framework is optimized using the AdamW optimizer [loshchilov2017decoupled] with a weight decay of $0.01$ and a polynomial learning rate decay schedule (power =0.9=0.9). Structurally, both students utilize a ViT-B/16 backbone [dosovitskiy2020image] and employ decoders with identical architectures but independent parameters. Regarding data augmentation, we strictly follow the protocols established in UniMatch [yang2023revisiting]. Specifically, labeled images and “weak” unlabeled views undergo basic geometric transformations, including random resizing (scale range 0.5-2.0) and horizontal flipping. “strong” unlabeled views are generated via an aggressive perturbation chain comprising color jittering, random grayscale conversion, Gaussian blur, and CutMix [yun2019cutmix]. Furthermore, feature-space perturbation is applied via dropout [srivastava2014dropout] (rate =0.5=0.5) within the dual-student forward mechanism. The unsupervised loss is guided by confidence-thresholded pseudo-labels (threshold =0.95=0.95). Regarding the training schedule, each experiment runs for 80 epochs, with the exception of the MER dataset [li2022stepwise]. Due to its limited sample size, MER requires a prolonged schedule of 150 epochs to mitigate noisy updates and ensure stable convergence. All experiments were conducted on a workstation equipped with NVIDIA RTX 3090 GPUs. We utilize the mean intersection-over-union (mIoU) as the evaluation metric.

### v2

@@ III-B Explicit-Implicit Semantic Collaborative Guidance
For the CLIP-based student, we adopt a concept-based prompt ensembling strategy to improve semantic precision. For each semantic category $k\in\{1,\dots,N\}$, we curate a set of fine-grained concept descriptions $\mathcal{C}_{k}=\{c_{k,i}\}_{i=1}^{S_{k}}$ (e.g., for road: {highway, main road, street, …}), as specified in the dataset configurations, where $S_{k}$ denotes the number of concepts contained in $\mathcal{C}_{k}$. Each concept $c_{k,i}\in\mathcal{C}_{k}$ is formatted by a prompt template function $\mathcal{T}(\cdot)$ (e.g., $\mathcal{T}(c_{k,i})=\text{``a photo of a \{concept\}''}$) and then encoded by the frozen CLIP text encoder $\mathcal{E}_{\text{txt}}$. We aggregate these concept embeddings by averaging to obtain a class prototype $\mathbf{t}_{k}\in\mathbb{R}^{C_{1}}$:
$\mathbf{t}_{k}=\frac{1}{S_{k}}\sum_{i=1}^{S_{k}}\mathcal{E}_{\text{txt}}\big(\mathcal{T}(c_{k,i})\big).$ (5)
Finally, we stack the prototypes of all $N$ categories into an explicit text-query matrix $\mathbf{F}_{\text{clip}}^{t}\in\mathbb{R}^{N\times C_{1}}$:
$\mathbf{F}_{\text{clip}}^{t}=[\mathbf{t}_{1},\dots,\mathbf{t}_{N}],$ (6)
@@
In contrast, for the DINOv3-based student, we utilize learnable queries $\mathbf{Q}\in\mathbb{R}^{N\times C_{2}}$ to provide implicit class-level guidance [56, 43]. These queries are implemented as a trainable embedding layer and are optimized alongside the DINOv3 backbone and the segmentation decoder:
$\mathbf{Q}=[\mathbf{q}^{1},\dots,\mathbf{q}^{N}].$ (7)
@@
Let $\mathbf{F}_{\text{clip}}^{\text{global}}\in\mathbb{R}^{C_{1}\times h\times w}$ and $\mathbf{F}_{\text{dino}}^{\text{local}}\in\mathbb{R}^{C_{2}\times h\times w}$ denote the visual features from the CLIP-based and DINOv3-based students, respectively. Both students employ an isomorphic yet parameter-independent query-based decoder, and compute segmentation logits by dot-product similarity:
$\mathbf{Z}^{\text{global}}=\mathrm{Sim}(\mathbf{F}_{\text{clip}}^{t},\mathbf{F}_{\text{clip}}^{\text{global}}),\quad\mathbf{Z}^{\text{local}}=\mathrm{Sim}(\mathbf{Q},\mathbf{F}_{\text{dino}}^{\text{local}}),$ (8)
where $\mathbf{Z}^{\text{global}},\mathbf{Z}^{\text{local}}\in\mathbb{R}^{N\times h\times w}$, and $\mathrm{Sim}(\cdot,\cdot)$ denotes a dot-product similarity implemented with an einsum operator. This formulation maintains a unified architectural structure while allowing the two students to leverage distinct semantic sources, namely explicit text embeddings versus implicit learnable queries.
@@ III-C Global-Local Feature Collaborative Fusion
Conventional co-training frameworks [4, 26, 15] typically rely on identical network architectures with different random initializations. Despite the divergent starting points, these homogeneous models often converge to similar error patterns due to confirmation bias. To mitigate this error accumulation, Co2S introduces a global-local feature collaborative fusion strategy. As visually evidenced in Fig. 3, the attention maps reveal distinct representational characteristics where CLIP produces diffuse activations rich in global contextual information while DINOv3 generates sharp responses highlighting local details. Building on this complementarity, our mechanism effectively fuses the global contextual information captured by CLIP [27] with the local details produced by DINOv3 [32]. This complementary fusion enables the model to generate segmentation maps with more accurate semantics and sharper, well-defined boundaries.
Given an unlabeled image $\mathbf{x}_{u}$, its weakly augmented view is fed into the two students to obtain the global logits $\mathbf{Z}^{\text{global}}\in\mathbb{R}^{N\times H\times W}$ and the local logits $\mathbf{Z}^{\text{local}}\in\mathbb{R}^{N\times H\times W}$. We first compute the confidence maps $\mathbf{C}_{\text{pred}}^{\text{global}}\in\mathbb{R}^{1\times H\times W}$ and $\mathbf{C}_{\text{pred}}^{\text{local}}\in\mathbb{R}^{1\times H\times W}$:
$\mathbf{C}_{\text{pred}}^{\text{global}}=\max_{n}\big[\mathrm{Softmax}(\mathbf{Z}^{\text{global}})\big]_{n},$ (9)
$\mathbf{C}_{\text{pred}}^{\text{local}}=\max_{n}\big[\mathrm{Softmax}(\mathbf{Z}^{\text{local}})\big]_{n},$ (10)
where $\mathrm{Softmax}(\cdot)$ is applied over the class dimension. Let $c_{\text{pred}}^{\text{global}}$ and $c_{\text{pred}}^{\text{local}}$ denote the scalar confidence values at a specific spatial location from $\mathbf{C}_{\text{pred}}^{\text{global}}$ and $\mathbf{C}_{\text{pred}}^{\text{local}}$, respectively. The joint maximum confidence for this pixel is defined as:
@@
We then build interaction masks based on three confidence scenarios: (1) If $c_{\text{pred}}^{\text{global}}\geq\tau$ and $c_{\text{pred}}^{\text{local}}\geq\tau$, the lower-confidence student learns from the higher-confidence one, when their confidence is equal, they engage in symmetric mutual learning. (2) If $C\geq\tau$ but one of $\{c_{\text{pred}}^{\text{global}},\,c_{\text{pred}}^{\text{local}}\}$ is below $\tau$, the low-confidence student is supervised by the high-confidence peer. (3) If $C<\tau$, the pixel is ignored and does not contribute to stabilization. Formally, the stability loss is defined as:
@@
where $\text{MSE}(\cdot,\cdot)$ represents the mean squared error calculation, $\mathbf{p}_{\ell},\mathbf{p}_{h}\in\mathbb{R}^{N}$ denote the class-probability vectors of the lower- and higher-confidence students at each pixel, respectively. They are computed by applying $\mathrm{Softmax}(\cdot)$ to the corresponding logits, $\tau$ is the confidence threshold, and $\lambda(t)$ is a ramp-up weight depending on the training step $t$.
@@ IV-A1 WHDLD
WHDLD [31] comprises 4,940 optical images captured by GaoFen-1 and ZiYuan-3 satellites, each having a size of $\times256$ pixels and a 2-meter spatial resolution. It provides pixel-wise annotations for six categories: building, road, pavement, vegetation, bare soil, and water.
@@ IV-A2 LoveDA
LoveDA [39] provides 5,987 images of size $\times1,024$ pixels with 0.3-meter resolution. Spanning urban and rural scenes, it annotates seven categories: background, building, road, water, barren, forest, and agriculture.
@@ IV-A3 Potsdam
Potsdam [30] consists of 38 aerial orthophotos with 5-cm resolution and $\times6,000$ dimensions. It covers six classes: impervious surface, building, low vegetation, tree, car, and clutter. These images are cropped into 5,472 non-overlapping $\times512$ patches for training and evaluation.
@@ IV-A5 MER and MSL
IV-B Implementation Details
The proposed Co2S framework is optimized using the AdamW optimizer [20] with a weight decay of $0.01$ and a polynomial learning rate decay schedule (power =0.9=0.9). Structurally, both students utilize a ViT-B/16 backbone [5] and employ decoders with identical architectures but independent parameters. Regarding data augmentation, we strictly follow the protocols established in UniMatch [47]. Specifically, labeled images and “weak” unlabeled views undergo basic geometric transformations, including random resizing (scale range 0.5-2.0) and horizontal flipping. “strong” unlabeled views are generated via an aggressive perturbation chain comprising color jittering, random grayscale conversion, Gaussian blur, and CutMix [51]. Furthermore, feature-space perturbation is applied via dropout [35] (rate =0.5=0.5) within the dual-student forward mechanism. The unsupervised loss is guided by confidence-thresholded pseudo-labels (threshold =0.95=0.95). Regarding the training schedule, each experiment runs for 80 epochs, with the exception of the MER dataset [18]. Due to its limited sample size, MER requires a prolonged schedule of 150 epochs to mitigate noisy updates and ensure stable convergence. All experiments were conducted on a workstation equipped with NVIDIA RTX 3090 GPUs. We utilize the mean intersection-over-union (mIoU) as the evaluation metric.
IV-C Comparison with State-of-the-Art Methods
We report comparison results across six widely used remote-sensing benchmarks to evaluate the performance of the proposed Co2S comprehensively. We conduct extensive comparisons against the supervised baseline (“OnlySup”), which is trained exclusively on the labeled data, as well as several general semi-supervised baselines, including FixMatch [33], UniMatch [47], U2PL [42], and remote-sensing-specific approaches like WSCL [21], DWL [12], and MUCA [41].
@@ IV-C Comparison with State-of-the-Art Methods
FixMatch [33]  [NeurIPS’20] 56.0 57.6 59.8 60.8
U2PL [42]  [CVPR’22] 57.1 58.3 59.9 61.1
WSCL [21]  [TGRS’23] 56.8 58.6 59.8 61.1
UniMatch [47]  [CVPR’23] 57.4 58.8 60.4 61.5
DWL [12]  [ISPRS’24] 56.5 57.0 57.1 58.9
MUCA [41]  [TGRS’25] 56.5 58.2 60.0 60.5
@@
TABLE II: Segmentation results (mIoU %) on the LoveDA dataset under different labeled ratios. The best and second-best results are highlighted in bold and underlined, respectively.
Method 1/40 1/16 1/8 1/4
OnlySup 45.9 53.1 56.6 58.9
FixMatch [33]  [NeurIPS’20] 53.3 57.9 60.1 61.5
U2PL [42]  [CVPR’22] 52.3 58.0 59.6 61.5
WSCL [21]  [TGRS’23] 54.1 57.4 59.4 61.1
UniMatch [47]  [CVPR’23] 55.5 58.6 60.0 61.7
DWL [12]  [ISPRS’24] 57.4 59.0 62.8 63.4
MUCA [41]  [TGRS’25] 54.6 56.7 60.2 63.1
Co2S (Ours) 58.2 60.4 62.7 64.0
@@
FixMatch [33]  [NeurIPS’20] 69.4 71.5 73.5 75.7
U2PL [42]  [CVPR’22] 67.2 71.3 73.8 75.9
WSCL [21]  [TGRS’23] 69.9 72.3 73.9 75.4
UniMatch [47]  [CVPR’23] 70.7 72.6 74.8 76.4
DWL [12]  [ISPRS’24] 74.2 78.3 79.8 80.3
MUCA [41]  [TGRS’25] 66.9 71.5 76.0 79.2
@@
TABLE IV: Segmentation results (mIoU %) on the MER and MSL datasets under different labeled ratios. The best and second-best results are highlighted in bold and underlined, respectively.
Method MER MSL
1/8 1/4 1/8 1/4
OnlySup 52.8 55.3 53.6 59.0
FixMatch [33]  [NeurIPS’20] 53.8 58.1 56.8 60.5
U2PL [42]  [CVPR’22] 54.9 56.1 55.5 61.6
WSCL [21]  [TGRS’23] 51.9 54.9 58.2 59.9
UniMatch [47]  [CVPR’23] 56.0 58.1 57.9 62.8
DWL [12]  [ISPRS’24] 54.5 54.7 59.0 63.6
MUCA [41]  [TGRS’25] 52.6 53.0 52.4 59.3
Co2S (Ours) 56.8 59.1 60.9 65.9
TABLE V: Segmentation results (mIoU %) on the GID-15 dataset under different labeled ratios. The best and second-best results are highlighted in bold and underlined, respectively.
Method OnlySup FixMatch [33] U2PL [42] WSCL [21] UniMatch [47] DWL [12] MUCA [41] Co2S (Ours)
@@
IV-C1 Quantitative Comparison and Analysis
The comprehensive comparisons across six diverse benchmarks (Tables I-V) demonstrate that Co2S consistently achieves leading performance, particularly in regimes with extremely scarce annotations. On the WHDLD [31] dataset, Co2S outperforms the strong competitor UniMatch by 3.7% under the extremely sparse 1/24 split, demonstrating its capability to parse diverse land-cover features accurately. Similarly, on the LoveDA [39] dataset, the method yields a remarkable 12.3% improvement over the supervised baseline at the 1/40 ratio, showing superior performance against significant urban-rural distribution shifts. For the ultra-high-resolution Potsdam [30] dataset, Co2S achieves the best performance at the 1/32 split with 74.3% mIoU, surpassing remote-sensing-specific designs like DWL. Extending to the large-scale GID-15 [38] dataset, our method outperforms UniMatch by 1.5% at the 1/8 split, indicating proficiency in handling diverse land-cover characteristics across vast spatial areas. Furthermore, in the extraterrestrial environments of MER and MSL [18], Co2S delivers the highest accuracy on these benchmarks characterized by extreme class imbalance. Crucially, a consistent trend observed across all benchmarks is that the performance advantage of Co2S over baselines amplifies as the labeled ratio decreases. This phenomenon validates the stability of our Co2S, confirming that the synergistic foundation model priors effectively prevent model collapse and sustain high accuracy even in the most extreme low-data annotation regimes.
IV-C2 Qualitative Visualization Analysis